### v1

# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

Abstract
Stably stratified shear flows, where the base velocity shear is quasi-continuously forced externally, can arise in many geophysically and environmentally relevant circumstances, for example at estuarine outflows. It is important to determine the emergent dynamics of the ensuing statistically steady stratified turbulence. We investigate this phenomenon in a series of three-dimensional direct numerical simulations using spectral element methods. We force the flow to relax back towards vertical hyperbolic tangent profiles of streamwise velocity and buoyancy, with characteristic half-depth $d_{0}$, half-velocity jump $U_{0}$, and half-buoyancy jump $B_{0}$, with a relaxation time $t_{r}=100\tau_{\textrm{adv}}$ where $\tau_{\textrm{adv}}:=d_{0}/U_{0}$. We consider computational domains of vertical extent $L_{z}=48$ (scaled with $d_{0}$) with a range of (scaled) horizontal extents $16\leq L_{\mathrm{h}}\leq 512$. We simulate a fluid with Prandtl number $\mathrm{Pr}:=\nu/\kappa=1$, where $\nu$ is the kinematic viscosity and $\kappa$ is the buoyancy diffusivity, and set the initial bulk Reynolds number $\mathrm{Re}_{0}:=U_{0}d_{0}/\nu=50$ and initial bulk Richardson number $\mathrm{Ri}_{0}:=B_{0}d_{0}/U_{0}^{2}=1/80$ (corresponding to the initial minimum gradient Richardson number). At these parameters, the flow is initially (vigorously) unstable to a primary Kelvin-Helmholtz instability. We simulate the continuously forced flow over about $5000\tau_{\textrm{adv}}$, and investigate the dynamically emergent length scales and turbulence properties of the statistically stationary flow, in particular the local turbulent flux coefficient $\Gamma_{\chi}$, as well as streamwise and spanwise Fourier spectra. We find that the shear layer half depth converges to $d\approx 8$ (i.e., $\Lambda_{z}\approx 16$) with markedly increased minimum gradient Richardson number ($\sim 0.1$) and $\mathrm{Re}\approx 400$ with associated convergent vertical mixing properties only for $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. However, emergent dominant yet large-scale spanwise or streamwise flow structures appear to extend up to (scaled) $\Lambda_{y}\approx 50$ or even $\Lambda_{x}\approx 115$, respectively, and thus may only be fully resolved for horizontal domains of extent $L_{\mathrm{h}}\gtrsim 256$. Our observations demonstrate the marked anisotropy of characteristic emergent length scales, even for such ‘weakly stratified’ forced shear flows, and are consistent with the possibility that an ‘imprint’ of the primary linear instability even continues to survive in such vigorously turbulent flows, as the streamwise flow structures have a scale consistent with the most unstable KHI for the converged deepened turbulent shear layer.
keywords:
forced stratified shear flow, emergent phenomena, anisotropy, pattern formation, direct numerical simulation
1 Introduction
Stably stratified (vertical) shear flows, where both the background fluid buoyancy (i.e. the appropriatedly scaled negative density perturbation) and horizontal flow velocity vary with height, are ubiquitous in geophysical contexts. There has been a very large body of work considering the ways in which such flows behave, (as evidenced by a sequence of reviews such as Fernando (1991); Peltier & Caulfield (2003); Ivey et al. (2008); Caulfield (2020) and Caulfield (2021)) with significant focus on the (possible) existence of turbulence extracting and converting the kinetic energy in the background shear and the associated enhanced (irreversible) turbulent mixing. Understanding (and parameterising) such turbulent mixing of stratified fluids is a key challenge in geophysical flows, as the transport of momentum, heat and other scalars (such as dissolved gases, pollution and microplastics for example) is both a crucial process and a phenomenon of great uncertainty for the description of the world’s climate and environment (see for example the reviews of Wunsch & Ferrari (2004); Ferrari & Wunsch (2009); Gregg et al. (2018)).
Figure 1: Fundamental configuration of forced stratified shear flows, with an initial buoyancy profile $b_{0}$ that is statically stable, and an imposed initial velocity profile $u_{x,0}$ which may induce flow instability leading to a transtion to turbulence. Right: A continuous forcing $f_{\Phi}$ injects the required energy to sustain this turbulence, ensuring statistically stationary dynamics.
In essence, as shown schematically in figure 1, stably stratified shear flows are characterised by a competition between a stabilising buoyancy and a de-stabilising velocity (or shear) profile. However, understanding fundamental aspects of this deceptively simple set-up is exceptionally complicated. Considering initial value problems of (initially) laminar velocity and density profiles, it is well-known that such flows can be prone to a range of primary flow instabilities (see the review of Caulfield (2021)) that effectively rearrange the strip of spanwise vorticity into trains of elliptical vortices (such as the classic Kelvin-Helmholtz ‘billows’, i.e. the saturated manifestation of the Kelvin-Helmholtz Instability (KHI)) which themselves are prone (at least for sufficiently high Reynolds number) to a ‘zoo’ of secondary instabilities (a nomenclature proposed by Mashayek & Peltier (2012a, b)) which trigger turbulence transition and hence significantly enhanced mixing and dissipation. As a (perhaps tenuous) paradigm for mixing induced by breaking internal waves in the atmosphere and ocean, direct numerical simulations of such transient mixing events have been widely conducted. There is then an inevitable set of trade-offs in how best to utilise finite computational resources. The need to simulate the flow at as high a Reynolds number as possible (defined in terms of the shear layer half depth $d_{0}$ and half velocity difference $U_{0}$ as shown in figure 1) to ensure sufficiently realistic turbulence must be balanced against the need for the computational domain to be sufficiently large to allow realistic flow structures, as well as the need to compute for sufficiently long times to capture appropriately the full life cycle of such mixing events.
Although such mixing events can extend for a significant length of time (particularly when the flow is prone primarily to the so-called ‘Holmboe Wave Instability’ (HWI) (Holmboe, 1962), as discussed in detail by Salehipour et al. (2016, 2018)), there will always be an inevitable decay, unless the underlying driving mechanism (the shear) is replenished. The particular life cycle of such mixing events can be strongly sensitive to initial conditions, as demonstrated conclusively by Liu et al. (2022), due to competition between different members of the secondary instability ‘zoo’, especially involving the subharmonic ‘pairing’ instability of neighbouring primary Kelvin-Helmholtz billows. As demonstrated by Mashayek & Peltier (2013), sufficiently strong stratification can disrupt and hence suppress such pairing (or more precisely ‘merging’) events due to the energetic costs of the inherent vertical motions. Such suppression is often cited as a reason to restrict computational domains considering such mixing events to one (or at most two) wavelengths of the primary instabilities, allowing the simulations to be conducted at sufficiently high Reynolds number for the ensuing mixing not to be (excessively) dominated by viscous and diffusive effects.
However, restricting the horizontal extent of the computational domain inevitably affects the flow dynamics also. As shown by Scinocca & Ford (2000), rich dynamics can occur in longer streamwise domains, where primary instabilities with close wavelengths (and linear growth rates) can compete as they grow to finite amplitude if the (conventional) imposition of periodic streamwise boundary conditions does not quantise the possible instabilities too severely. Analogous issues arise with the spanwise extent of the computational extent. Many studies consider relatively narrow computational domains, in the sense that the spanwise extent is (often significantly) smaller than the characteristic (streamwise) wavelength of the primary instability. Such domains allow many of the (essentially local) secondary instabilities to develop and hence trigger turbulence transition. However, as can be straightforwardly observed in sufficiently wide tilting tank experiments (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and sufficiently wide computational domains, as clearly demonstrated by Fritts et al. (2022a, b), inherently three-dimensional ‘knots’, ‘tubes’, and billow ‘defects’ can develop in the spanwise direction on scales of the order (but typically larger) of the primary instability’s (streamwise) wavelength.
Although such initial-value-problem mixing events are undoubtedly of geophysical interest and indeed demonstrate that the chosen size of the computational domain can affect the emergent flow structures, it is exceptionally difficult to determine to what extent the size of the considered flow domain affects the ensuing stratified turbulence (and associated quantities of interest, such as the mixing), due not least to the inherent transient nature of the mixing event. However, it is important to remember that such initial-value-problem mixing events are really just one end member class of the possible flows of geophysical interest.
The other obvious end member class is the class of continuously forced flows, where the ‘background’ velocity and density (or equivalently buoyancy) profiles are driven by some external, quasi-steady forcing. Possible candidate mechanisms for such geophysically relevant forcings include wind, solar radiation and resulting evaporation at the surface of the ocean (Thorpe, 2005), tidal forcing (Laurent et al., 2002), active matter living in water (Castro et al., 2022), or continuous outflows from rivers (Uncles & Mitchell, 2011). While this list is not meant to be complete in any sense, it illustrates that quasi-steady forcings do occur in geophysically relevant situations. An (artificial) volumetric forcing is a particularly convenient way to mimic these natural complex mechanisms in a simplified manner. As visualised on the right side of figure 1, such a forcing may be defined to relax the local profiles.
Such a forcing was used by Smith et al. (2021), who demonstrated that, after an initial transient where primary instabilities (either KHI or HWI) develop and (inevitably) break down, the ensuing turbulence could be sustained over arbitrarily long times, thus enabling statistically steady statistics of the flow to be calculated. There is a clear attraction to considering such forced flows due principally to the inherent removal of the confounding effects of the life cycle of the mixing event from the turbulence statistics. Therefore, such flows seem the natural test-bed to consider the effect of the size of the computational domain on the convergence (or not) of such statistics, in an attempt to identify how large a computational domain needs to be for the imposition of horizontal periodic boundary conditions to become insignificant. However, as in the initial value problem flows discussed above, Smith et al. (2021) also considered relatively small computational domains (in terms of the initial shear layer half-depth $d_{0}$), and so it is (again) not at all clear whether or not the flow remained unaffected by the computational domain size. An equivalent question to ask is what are the (unconstrained) emergent length scales of a forced stratified shear flow, and answering that (open) question is the central objective of this paper.
To address this question, the rest of the paper is organised as follows. In section 2 we present our numerical approach before we list the range of considered computational domains in section 3. For simplicity, we focus on a relatively weakly stratified flow, prone to primary KHI, that initially ‘rolls up’ into a train of Kelvin-Helmholtz billows. We identify some interesting early-time dynamics that may possibly be affected by diffusion. However, it is not this state that is of principal interest, but rather the long-time statistically steady (yet sheared) turbulent state. We present the results of those simulations, identifying the (strongly anisotropic) length scales which emerge in sufficiently large flow domains. We also demonstrate conclusively that key statistics of the flow, including in particular those related to irreversible mixing, are sensitive to the size of the computational domain. We remark that for convergent statistics, perhaps surprisingly, the flow domain needs to be extraordinarily ‘large’ or extended compared to the initial shear-layer half-depth (i.e. of order $100$ times larger). Finally, in section 4 we draw our conclusions, in particular discussing the implications of our results for future research.
2 Numerical approach
2.1 Governing equations
We consider an incompressible flow based on the Oberbeck-Boussinesq approximation with a linear equation of state. The three-dimensional equations of motion are non-dimensionalised using the (dimensional) initial magnitudes of the streamwise velocity $U_{0}$ and buoyancy $B_{0}$, as well as the shear layer half-depth $d_{0}$ as shown in figure 1. Using the advective time scale $\tau_{\textrm{adv}}:=d_{0}/U_{0}$ and the appropriate characteristic pressure scale $p_{\textrm{char}}:=U_{0}^{2}\rho_{\textrm{ref}}$, non-dimensional variables (marked here with a tilde) can be related to the dimensional variables as follows:
$\bm{x}=d_{0}\tilde{\bm{x}},\qquad\bm{u}=U_{0}\tilde{\bm{u}},\qquad b=B_{0}\tilde{b},\qquad t=\tau_{\textrm{adv}}\tilde{t},\qquad p=p_{\textrm{char}}\tilde{p}$ (1)
Henceforth, we focus on non-dimensional variables, and so drop the tildes from all variables.
As a consequence, the non-dimensional governing equations are
$\nabla\cdot\bm{u}$ $=0,$ (2)
$\frac{\partial\bm{u}}{\partial t}+\left(\bm{u}\cdot\nabla\right)\bm{u}$ $=-\nabla p+\frac{1}{\mathrm{Re}_{0}}\nabla^{2}\bm{u}+\mathrm{Ri}_{0}b\bm{e}_{z}+f_{u}\bm{e}_{x},$ (3)
$\frac{\partial b}{\partial t}+\left(\bm{u}\cdot\nabla\right)b$ $=\frac{1}{\mathrm{Re}_{0}\mathrm{Pr}}\nabla^{2}b+f_{b},$ (4)
where $\bm{u}$, $b$ and $p$ represent the velocity, buoyancy, and modified pressure field, respectively. The precise form of the volumetric forcing terms $f_{u}$ and $f_{b}$ will be defined below. The buoyancy $b:=-\rho^{{}^{\prime}}g/\rho_{\textrm{ref}}$ and corresponds to the negative of the reduced gravity, so $\rho^{\prime}$ is the deviation from the reference density $\rho_{\textrm{ref}}$. Three non-dimensional parameters naturally emerge from this scaling: the Prandtl number, the initial bulk Reynolds number, and the initial bulk Richardson number,
$\mathrm{Pr}:=\frac{\nu}{\kappa},\qquad\mathrm{Re}_{0}:=\frac{U_{0}d_{0}}{\nu},\qquad\mathrm{Ri}_{0}:=\frac{B_{0}d_{0}}{U_{0}^{2}},$ (5)
respectively, where $\nu$ is the kinematic viscosity and $\kappa$ is the buoyancy diffusivity. We restrict attention to statically stable situations where $\mathrm{Ri}_{0}>0$.
The volumetric forcing terms $f_{u}$ and $f_{b}$ in equations (3) and (4) are a crucial aspect of our configuration. Following Smith et al. (2021), we consider
$f_{u}$ $:=-\frac{1}{t_{\textrm{r}}}\left[u_{x}-u_{x,0}\right]$ with $u_{x,0}\left(z\right)$ $:=\tanh\left(z\right),$ (6)
$f_{b}$ $:=-\frac{1}{t_{\textrm{r}}}\left[b-b_{0}\right]$ with $b_{0}\left(z\right)$ $:=\tanh\left(R_{0}z\right)$ (7)
where $t_{\textrm{r}}$ is the response time while $u_{x,0}$ and $b_{0}$ represent the initial streamwise velocity and buoyancy base profiles to which the flow relaxes back. In this context, $R_{0}:=d_{0}/\delta_{0}$ defines the ratio of initial interface half-depths (i.e. $\delta_{0}$ represents the dimensional initial buoyancy interface half-depth) with $R_{0}=\sqrt{\mathrm{Pr}}$ following the diffusive arguments presented by Smyth et al. (1988). In essence, these forcing terms are idealized models of geophysically relevant processes that tend to restore the initial profiles $u_{x,0}$ and $b_{0}$, and thus sustain turbulence over arbitrarily long times.
Hence, the governing equations (2) – (4) are fully specified via four control parameters: $\mathrm{Pr}$, $\mathrm{Re}_{0}$, $\mathrm{Ri}_{0}$, and $t_{\textrm{r}}$. The associated energy equations are provided in appendix A.
2.2 Domain, boundary and initial conditions, and numerical code
As indicated by figure 1, the streamwise, spanwise, and vertical extents of the Cartesian numerical domain are $L_{x}$, $L_{y}$, and $L_{z}$, respectively. Both the midpoint of the shear layer and the midpoint of the buoyancy distribution are located at midplane, $z=0$, with a horizontal cross-section $A:=L_{x}\times L_{y}$. We consider a horizontally periodic domain where any quantity
$\Phi\left(\bm{x}\right)=\Phi\left(\bm{x}+i_{x}L_{x}\bm{e}_{x}+i_{y}L_{y}\bm{e}_{y}\right)\qquad\textrm{given}\quad i_{x,y}\in\mathbb{N}.$ (8)
Additionally, we apply free-slip and no-flux boundary conditions at the top and bottom, so
$u_{z}=\frac{\partial u_{x}}{\partial z}=\frac{\partial u_{y}}{\partial z}=0\qquad\textrm{and}\qquad\frac{\partial b}{\partial z}=0\qquad\textrm{at}\quad z=\pm\frac{L_{z}}{2}.$ (9)
Our initial condition is given by
$u_{x}=u_{x,0},\qquad u_{y}=u_{z}=0,\qquad\textrm{and}\qquad b=b_{0}+\Upsilon\qquad\textrm{at}\quad t=0.$ (10)
The random fluctuations $-10^{-3}\leq\Upsilon\left(\bm{x}\right)\leq 10^{-3}$ ‘seed’ the development of primary instabilities.
We solve the coupled governing equations (2) – (4), subject to these boundary and initial conditions, using the GPU-accelerated spectral element solver NekRS (Fischer, 1997; Scheel et al., 2013; Fischer et al., 2022). As shown in more detail in appendix B, spectral element methods can accommodate different required spatial resolutions across the domain and are thus perfectly suited to resolve shear flows efficiently. This is particularly important given degrees of freedom of up to $N_{\textrm{dof}}\approx 3N_{\textrm{e}}N^{3}\sim\mathcal{O}\left(10^{9}\right)$ in our present study.
$L_{x}\times L_{y}\thickspace\thickspace$ $N_{\mathrm{e},x}\times N_{\mathrm{e},y}$ $d$ $\delta$ $R$ $\mathrm{Re}$ $\mathrm{Ri}$
$16^{2}$ 00 $9^{2}$ 00 $2.66\pm 0.16$ $2.80\pm 0.14$ $0.95\pm 0.04$ $133\pm\hphantom{0}8$ $0.0333\pm 0.0020$
$32^{2}$ 00 $18^{2}$ 00 $5.35\pm 0.34$ $5.54\pm 0.35$ $0.96\pm 0.02$ $267\pm 17$ $0.0668\pm 0.0043$
$64^{2}$ 00 $36^{2}$ 00 $7.55\pm 0.47$ $7.75\pm 0.36$ $0.97\pm 0.03$ $377\pm 24$ $0.0943\pm 0.0059$
$96^{2}$ 00 $54^{2}$ 00 $7.92\pm 0.53$ $8.27\pm 0.56$ $0.96\pm 0.02$ $396\pm 26$ $0.0990\pm 0.0066$
$128^{2}$ 00 $72^{2}$ 00 $8.05\pm 0.42$ $8.41\pm 0.45$ $0.96\pm 0.01$ $403\pm 21$ $0.1007\pm 0.0053$
$256^{2}$ 00 $144^{2}$ 00 $7.90\pm 0.18$ $8.22\pm 0.17$ $0.96\pm 0.01$ $395\pm\hphantom{0}9$ $0.0987\pm 0.0022$
$512^{2}$ 00 $288^{2}$ 00 $7.73\pm 0.08$ $8.06\pm 0.07$ $0.96\pm 0.00$ $387\pm\hphantom{0}4$ $0.0967\pm 0.0010$
$2048\times 512$ $1152\times 288$ $7.79\pm 0.03$ $8.12\pm 0.03$ $0.96\pm 0.00$ $390\pm\hphantom{0}1$ $0.0974\pm 0.0003$
$512\times 2048$ $288\times 1152$ $7.77\pm 0.04$ $8.10\pm 0.04$ $0.96\pm 0.00$ $388\pm\hphantom{0}2$ $0.0971\pm 0.0005$
Table 1: Simulation parameters. The Prandtl number $\mathrm{Pr}=1$, initial bulk Reynolds number $\mathrm{Re}_{0}=50$, initial bulk Richardson number $\mathrm{Ri}_{0}=0.0125$, response time $t_{\textrm{r}}=100$, and initial ratio of interface (half) thicknesses $R_{0}=1$ in a horizontally periodic domain. In the vertical direction, the domain has an aspect ratio $L_{z}=48$ spanned by $N_{\mathrm{e},z}=18$ non-uniformly distributed spectral elements (see appendix B for more details) together with free-slip and no-flux boundary conditions for the velocity and buoyancy field, respectively. The polynomial order $N=6$. Although the total evolution or run time of each flow $t_{\textrm{evo}}=5,040$, this work focuses on the statistically stationary dynamics during the last $\Delta t=3,000$ only. For each simulation, this table lists the horizontal aspect ratios $L_{x}\times L_{y}$ and corresponding numbers of uniformly distributed spectral elements $N_{\mathrm{e},x}\times N_{\mathrm{e},y}$. Moreover, we include the final and dynamically manifesting shear layer (half) depth $d$ of the streamwise velocity field as well as the buoyancy interface half-depth $\delta$, their ratio $R$, the bulk Reynolds number $\mathrm{Re}$, as well as the bulk Richardson number $\mathrm{Ri}$, listing both temporal means and standard deviations.
3 Results
This study considers forced stratified shear flows at $\mathrm{Pr}=1$, $\mathrm{Re}_{0}=50$, $\mathrm{Ri}_{0}=0.0125$, and $t_{\textrm{r}}=100$ in domains of $L_{z}=48$. These parameters, as will be shown in more detail below, are associated with the minimum value of the initial gradient Richardson number $\mathrm{Ri}_{\textrm{g,0}}(z=0)=\mathrm{Ri}_{0}R_{0}$, which is sufficiently small in principle to allow the development of primary KHI. Note that the initial gradient Richardson number $\mathrm{Ri}_{\textrm{g,0}}(z)$ is defined in terms of the initial profile,
$\mathrm{Ri}_{\textrm{g,0}}(z):=\mathrm{Ri}_{0}\frac{\frac{\partial b_{0}}{\partial z}}{\left(\frac{\partial u_{x,0}}{\partial z}\right)^{2}}.$ (11)
We investigate and quantify the emergent dynamics of the flow while varying the horizontal extents of the domain, $L_{x}$ and $L_{y}$. Normally, we consider domains of square horizontal cross-section with $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ ranging from $16$ to $512$. Table 1 summarises all our considered domains.
3.1 Typical evolution of a forced stratified shear flow
Figure 2: Temporal evolution of forced, stratified shear flows. (a – e) The flow is prone to a primary KHI, leading eventually to ‘overturning billows’ and streamwise mergers. A continuous forcing sustains the induced turbulence for arbitrarily long times. During this evolution of the flow, (f – h) the interface broadens before reaching a statistically stationary depth. Note that for this relatively small $\mathrm{Re}_{0}$, as shown in (g), diffusion of the shear layer and density interface dominates the development of the primary instability until the shear layer and density interface have approximately doubled in depth. In this figure, $L_{\mathrm{h}}=128$ while panels (a – e) visualise $b\left(x,y=0,z,t\right)$ with the colour bar matching figure 6 (l, p).
We first consider the evolution of a typical flow in a domain with horizontal extent of $L_{\mathrm{h}}=128$. Figures 2(a – e) show snapshots of the instantaneous buoyancy field $b\left(x,y=0,z,t\right)$ in vertical slices normal to the spanwise direction. At early times, the flow is prone to a primary Kelvin-Helmholtz instability which, driven by the vertical shear, develops into a train of KH billows (panel (c)) that merge subsequently (panel (d)). These primary billows break down, and the buoyancy interface broadens (panel (f)). Note that we use $\langle\cdot\rangle_{\Phi}$ to denote averages over $\Phi$.
If this was an initial value problem, turbulence would die out shortly after $t=270$ due to the combined (and inter-related) effects of enhanced dissipation and broadening of both the shear layer as well as the density interface, eventually leading to an increased (and, according to the so-called Miles-Howard criterion (Miles, 1961; Howard, 1961), indeed linearly stable) minimum gradient Richardson number. However, volumetric forcing sustains the induced turbulence over arbitrarily long times. This is underlined by panel (e), which shows a snapshot during this late statistically steady turbulent state of the flow. Here, we focus largely on this late, statistically stationary dynamics.
Across the evolution of the flow, we quantify the associated broadening of the shear layer half-depth and buoyancy interface half-thickness via
[…]
respectively, using the approach proposed by (Smyth & Moum, 2000). Together with the initial profiles defined in equations (6) and (7), these functions yield $d\left(t=0\right)=1$ and $\delta\left(t=0\right)=1/R_{0}$ (as these non-dimensional lengths $\left\{d,\delta\right\}$ are measured in units of $d_{0}$, see again equation (1)). Similarly, the time-dependent ratio of interface (half) thicknesses $R\left(t\right):=d/\delta$ with $R\left(t=0\right)=R_{0}$. Figures 2 (g, h) highlight that both $d$ and $\delta$ converge after an initial transient to statistically stationary values $\left\{d,\delta\right\}\gg 1$. In other words, the interfaces have deepened significantly. Their ratio $R\approx 1$ at late times, consistently with the fact that the flow remains ‘weakly’ stratified. Here, ‘weakly stratified’ is meant in the specific sense that the turbulent diffusivity of buoyancy closely follows the turbulent diffusivity of the momentum. Equivalently, the turbulent Prandtl number is close to one, and so the buoyancy field is at least in some sense slaved to the velocity field and behaves like a passive scalar. We note in passing that, as highlighted in panel (g), the onset of the primary KHI (evidenced by the significant change in the rate of the growth of $d$ and $\delta$ around $t\simeq 150$) occurs only after a period of significant diffusive deepening of both the shear layer and the buoyancy interface. As the shear layer depth $d$ approximately doubles, the effective Reynolds number (see below) approximately doubles as well, thus eventually triggering the onset of the primary instability with a significantly larger wavelength than that expected when defined in terms of the initial shear layer half-depth $d_{0}$. Investigation of this interesting early-time interaction between diffusion and instability onset is beyond the scope of this paper but undoubtedly worthy of further, more detailed consideration.
Moreover, the above definition of $d\left(t\right)$ enables the computation of time-dependent values of the bulk Reynolds and Richardson number via
$\mathrm{Re}\left(t\right):=\frac{U_{0}\left(d\thickspace d_{0}\right)}{\nu}=\mathrm{Re}_{0}\thickspace d\qquad\textrm{and}\qquad\mathrm{Ri}\left(t\right):=\frac{B_{0}\left(d\thickspace d_{0}\right)}{U_{0}^{2}}=\mathrm{Ri}_{0}\thickspace d$ (13)
with, again, $\mathrm{Re}\left(t=0\right)=\mathrm{Re}_{0}$ and $\mathrm{Ri}\left(t=0\right)=\mathrm{Ri}_{0}$. We stress that, since $\langle d\rangle_{t}\approx 8\gg 1$ at our late times of interest, the final flow has an effective temporal average $\langle\mathrm{Re}\rangle_{t}\approx 400\gg\mathrm{Re}_{0}$ and has, thus, become much more turbulent than what one might have expected from the relatively small initial value of $\mathrm{Re}_{0}=50$.
Table 1 lists the temporal averages and standard deviations of $d$, $\delta$, $R$, $\mathrm{Re}$, and $\mathrm{Ri}$ during the late statistically stationary dynamics of all our simulations. We discuss their trends with respect to the horizontal domain size $L_{\mathrm{h}}$ in section 3.3.
3.2 Dependence of horizontally-averaged fields on domain size
Figure 3: Evolution of the horizontally-averaged buoyancy field. (a) The deepening of the density interface is affected by the horizontal extent of the domain $L_{\mathrm{h}}\equiv L_{x}=L_{y}$ and converges at large values of $L_{\mathrm{h}}$ only. For $L_{\mathrm{h}}=128$: (b) the late temporal average of $\mathrm{Ri}_{\textrm{g}}$ shows a significant increase in its minimum value to $0.74$, with an associated increase throughout the turbulent ‘mixing zone’ (the canonical marginal value of $0.25$ is marked with a vertical dot-dashed line), while the resulting associated mixing in the mixing zone is underlined by high amplitudes in (c) the stabilising vertical buoyancy advection (i.e. the buoyancy flux) $B$ and (d) the dissipation rates of kinetic energy and scaled buoyancy variance, $\varepsilon_{u}$ and $\chi$.
Turning attention to domain size effects, in figure 3 (a) we plot the mean vertical buoyancy profiles $\langle b\rangle_{A,t}$ associated with the late statistically stationary dynamics. We find that these profiles depend strongly on the horizontal extent $L_{\mathrm{h}}$ of the domain. Although $L_{\mathrm{h}}=16\gg 1$ even for our smallest domain, the interface deepens for increasing $L_{\mathrm{h}}$ and a convergence seems to be reached for $L_{\mathrm{h}}\gtrsim 64$ only. This suggests that the vertical mixing of buoyancy strongly depends on the horizontal extent of the domain.
This observed deepening of the buoyancy interface is accompanied by a deepening of the shear layer, both affecting in turn the (time-dependent) gradient Richardson number
$\mathrm{Ri}_{\textrm{g}}\left(z,t\right):=\frac{\langle N^{2}\rangle_{A}}{\langle S\rangle_{A}^{2}}$ (14)
via the vertical shear $S=\partial u_{x}/\partial z$ and buoyancy frequency
$N:=\sqrt{\mathrm{Ri}_{0}\frac{\partial b}{\partial z}}.$ (15)
We remark that $\mathrm{Ri}_{\textrm{g}}\left(z,t=0\right)=\mathrm{Ri}_{\textrm{g,0}}$ as defined in (11). As shown in figure 3(b), the minimum value of $\mathrm{Ri}_{\textrm{g}}$ has significantly increased (to $\langle\mathrm{Ri}_{\textrm{g}}\left(z=0\right)\rangle_{t}\approx 0.074$ for $L_{\mathrm{h}}=128$). Note that, due to the square in the denominator of the definition, the reduction of shear overpowers the simultaneous reduction of buoyancy stratification.
From this panel (b), it is also apparent that $\mathrm{Ri}_{\textrm{g}}\leq 0.25$ within the region $-\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}$. Although the flow is undoubtedly turbulent – and so the Miles-Howard criterion (Miles, 1961; Howard, 1961) of (steady) linear inviscid stability theory is definitely not applicable –, such relatively small values of $\mathrm{Ri}_{\textrm{g}}$ are necessary for turbulence to survive and so we refer to this zone as the turbulent mixing zone.
Indeed, this nomenclature is also strongly justified (as shown in figures 3(c, d)) by considering the vertical buoyancy advection (frequently called the buoyancy flux)
$B:=\mathrm{Ri}_{0}u_{z}b$ (16)
as well as the dissipation rates of kinetic energy and scaled buoyancy variance
$\varepsilon_{u}$ $:=\frac{2}{\mathrm{Re}_{0}}\bm{S}^{2}\qquad$ $\textrm{with}\quad\bm{S}$ $:=\frac{1}{2}\left[\left(\nabla\bm{u}\right)+\left(\nabla\bm{u}\right)^{T}\right]\quad\textrm{and}$ (17)
$\chi$ $:=\frac{\mathrm{Ri}_{0}^{2}\thickspace\varepsilon_{b}}{\langle N^{2}\rangle_{A,t}}\qquad$ $\textrm{with}\quad\varepsilon_{b}$ $:=\frac{1}{\mathrm{Re}_{0}\mathrm{Pr}}\left(\nabla b\right)^{2},$ (18)
respectively. These (non-dimensional) quantities are related to the evolution equations of kinetic energy and buoyancy variance as shown in appendix A. Here, $\bm{S}$ represents the strain rate tensor and $\varepsilon_{b}$ the (unscaled) buoyancy variance dissipation rate. Note that the scaling via $\langle N^{2}\rangle_{A,t}$ results in identical physical units for the associated dimensional dissipation rates. We emphasise this point by the comparison
$\varepsilon_{u}$ $=\frac{U_{0}^{3}}{d_{0}}\thickspace\tilde{\varepsilon}_{u},$ (19)
$\varepsilon_{b}$ $=\frac{B_{0}^{2}U_{0}}{d_{0}}\thickspace\tilde{\varepsilon}_{b}$ $=\frac{U_{0}^{5}}{d_{0}^{3}}\mathrm{Ri}_{0}^{2}\thickspace\tilde{\varepsilon}_{b},$ (20)
$\chi$ $=\frac{B_{0}^{2}d_{0}}{U_{0}}\thickspace\tilde{\tilde{\chi}}$ $=\frac{U_{0}^{3}}{d_{0}}\mathrm{Ri}_{0}^{2}\thickspace\tilde{\tilde{\chi}}$ $=\frac{U_{0}^{3}}{d_{0}}\thickspace\tilde{\chi},$ (21)
where we have, for improved clarity, re-introduced tildes for non-dimensional quantities in the above three equations only. This property of identical physical units is particularly helpful when studying the exchange of kinetic and potential energy, as buoyancy variance is closely related to the concept of ‘available potential energy’ (Caulfield, 2021).
As shown in panels (c – d) of figure 3, all of these quantities exhibit pronounced peaks close to the midplane. However, while the vertical buoyancy advection $B$ introduces a macroscopic stirring that is generally reversible, due to the inherent coupling of $\bm{u}$ and $b$ it leads to irreversible dissipation via both $\varepsilon_{u}$ and $\varepsilon_{b}$ (Caulfield, 2021), see again appendix A. As $\langle B\rangle_{V,t}<0$ (with the domain volume $V:=A\times L_{z}$), this stirring comes with an overall stabilising effect on the configuration. Furthermore, although enhanced values of $\varepsilon_{u}$ and $\chi$ extend beyond the ‘mixing zone’, it is clear that the mixing zone contains the vast majority of the enhanced dissipation at these late times.
3.3 Convergence of vertical stirring, dissipation and mixing for extended domains
Figure 4: Impact of the horizontal extent of the domain on the mixing. The flow topology – comprising the (a) interface depths, (b) final emergent (bulk) Reynolds number and Richardson number, as well as averages across the midplane and across the entire mixing zone of (c) gradient Richardson number, (d) buoyancy flux, (e) kinetic energy dissipation and (f) buoyancy variance dissipation – only converges for horizontally extended domains, $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. Vertical solid lines indicate the temporal standard deviation.
After having introduced important quantities related to the mixing of the buoyancy field in the previous sections 3.1 and 3.2, in figure 4 we plot their variation as a function of $L_{\mathrm{h}}$. Underlining the conclusions from figure 3 (a), figure 4(a) shows that the (half-) thicknesses of the shear layer or the streamwise velocity and buoyancy interfaces converge for extended horizontal domains only. As the parameters $\left\{\mathrm{Re},\mathrm{Ri}\right\}\propto d$, they also only converge for extended domains, as shown in panel (b). Note both are significantly increased (effectively by a factor of eight) from their initial values. In panel (c) we also plot temporal averages of the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$ as defined in equation (14). Both our analyses for the midplane and the entire mixing zone demonstrate again a convergence for extended domains only. Unsurprisingly, we observe a similar behaviour for the mixing-related quantities $B$, $\varepsilon_{u}$, and $\chi$ as shown in panels (d – f).
In summary, we find that the properties of the (predominantly vertical) stirring, dissipation and mixing actually depend strongly on the horizontal extent of the domain $L_{\mathrm{h}}$. The associated emergent vertical dynamics of the flow seems to become independent of $L_{\mathrm{h}}$ only once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$ where $L_{\mathrm{h,crit}}$ is a critical extent of the domain.
Figure 5: Statistics of mixing properties at midplane. Statistical distributions of (a) kinetic energy dissipation rate $\varepsilon_{u}$, (b) scaled buoyancy variance dissipation rate $\chi$, and (c) local flux coefficient $\Gamma_{\chi}$ are affected by insufficient horizontal extents of the domain. The grey dashed vertical line marks the canonical value $\Gamma_{\chi}=0.2$. Note the emergent scaling for $\Gamma_{\chi}$ for extreme mixing events, and the marked difference of the high tails of the PDFs of $\varepsilon_{u}$ and $\chi$.
As shown in figure 5, the dependence of dissipation and mixing rates on $L_{\mathrm{h}}$ does not just affect mean values but even the general structure of their probability density functions (PDFs). Interestingly, we observe a pronounced scaling of the PDF of the local flux coefficient $\Gamma_{\chi}$ – defined as
$\Gamma_{\chi}:=\frac{\chi}{\varepsilon_{u}},$ (22)
and representing the local ratio between dissipation of scaled buoyancy variance and kinetic energy – for values $\Gamma_{\chi}\gtrsim 0.2$, i.e., the canonical value proposed by Osborn (1980), as indicated in panel (c).
Indeed, as is apparent from figure 4 (e, f), the converged values of ‘bulk’ turbulent mixing coefficients
$\bar{\Gamma}_{\chi,\Phi}:=\frac{\langle\chi\rangle_{\Phi,t}}{\langle\varepsilon_{u}\rangle_{\Phi,t}},$ (23)
where the average is taken across either $z=0$ or $-\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}$, are substantially smaller than the canonical maximum value of $0.2$. We find that $\bar{\Gamma}_{\Phi}\simeq 0.079-0.081$ at the midplane and $\bar{\Gamma}_{\Phi}\simeq 0.13-0.14$ within the mixing zone, essentially because $\varepsilon_{u}$ is more localized at the midplane than $\chi$ is. This discrepancy relative to the proposal of Osborn (1980) is not entirely surprising, as it is clear that though both $\varepsilon_{u}$ and $\chi$ are elevated in the mixing zone, the assumption that they are linearly related (which is at the heart of Osborn’s influential arguments) is clearly not qualitatively or quantitatively correct.
3.4 Anisotropy of emergent large-scale dynamics
While the previous section 3.3 has made clear that vertical aspects of the flow dynamics converge for horizontally highly extended domains only, section 3.1 has demonstrated that the forced shear flows considered here develop certain characteristic horizontal structures, such as the primary Kelvin-Helmholtz billows. Therefore, it is appropriate to investigate whether there are emergent horizontally-aligned structures in the statistically steady flow at later times (when such early billows have broken down).
Figure 6: Emergent horizontally extended dynamics. From (a – h) early to (i – p) late times, the size of emergent large-scale flow structures exhibits a strong anisotropy between their (a – d, i – l) horizontal and (e – h, m – p) vertical extension. This figure shows data from the largest square domain, $L_{\mathrm{h}}=512$, with $z=0$ in (a – d, i – l) or $y=0$ in (e – h, m – p). While the distinct streamwise ‘overturning billows’ and spanwise ‘knots’ and ‘tubes’ mentioned in the introduction are prominent at early times, these structures (at least superficially) disappear at later times.
Figure 6 visualises both the early and final emergent dynamics present in our largest square domain of $L_{\mathrm{h}}=512$. Panels (a – d, i – l) depict the entire horizontal cross-section at midplane, $\Phi\left(z=0\right)$, whereas panels (e – h, m – p) depict an associated vertical slice at $\Phi\left(y=0\right)$. Note that the vertical slices of $b$ from panels (h, p) remind of figure 2 (c, e) despite the domain now being $16$ times as large. A video of the evolution of the flow in this domain of $L_{\mathrm{h}}=512$ – from $t=0$ to $t=t_{\textrm{evo}}$ – is provided as supplemental material.
At the earlier time, as shown in panel (h), the flow is clearly associated with the growth of the initial KHI and the subsequent formation of overturning billows, analogously to the simulation discussed in section 3.1 and shown in figure 2. Interestingly, as becomes clear via the associated horizontal slice in panel (d), although the initial overturning billows may extend across the entire extended spanwise direction, knots, tubes and defects between these rolls may introduce defects and imperfections to this otherwise regular pattern, as previously observed experimentally (Thorpe, 1985, 1987; Caulfield et al., 1996; Thorpe, 2002) and numerically (Fritts et al., 2022a, b) for (in the spanwise direction) sufficiently wide domains. However, these early aspects of the flow dynamics appear (at least superficially) to be absent or smeared out by the sustained turbulence at late times as shown in panels (l, p).
As the velocity field, in particular $u_{x}$ as shown in panels (a, e, i, m), exhibits structures that are very similar to the buoyancy field $b$ shown in panels (d, h, l, p), the observations made in the previous sections are further supported. However, we find that both the initial pattern formation as well as the late sustained turbulence within the mixing zone can be recognised more easily from the scalar buoyancy field than from the vectorial velocity field.
Independently of the specific point in time, this comparison of vertical slices with the corresponding horizontal midplanes demonstrates the co-existence of apparently quasi-regular dynamics in both the streamwise and spanwise directions. However, these visualisations suggest that the horizontal structures have characteristic length scales much larger than the shear layer (half-)depth, i.e., $\left\{d,\delta\right\}\ll\Lambda_{\textrm{hor}}$, implying a strong scale separation or anisotropy in the flow.
As a first step towards a quantitative consideration of the streamwise and spanwise extension of the emergent horizontal dynamics of the flow, we compute the Fourier (energy or co-) spectra
$E_{\Phi_{1}\Phi_{2}}\left(k_{x},y,z=0,t\right)$ $:=C\thickspace\real\left(\hat{\Phi}_{1}\hat{\Phi}_{2}^{*}\right)\qquad$ $\textrm{with}\quad\hat{\Phi}_{1,2}\equiv\hat{\Phi}_{1,2}\left(k_{x},y,z=0,t\right),$ (24)
$E_{\Phi_{1}\Phi_{2}}\left(x,k_{y},z=0,t\right)$ $:=C\thickspace\real\left(\hat{\Phi}_{1}\hat{\Phi}_{2}^{*}\right)\qquad$ $\textrm{with}\quad\hat{\Phi}_{1,2}\equiv\hat{\Phi}_{1,2}\left(x,k_{y},z=0,t\right)$ (25)
of various quantities. Here, $\hat{\Phi}$ represent the Fourier coefficients corresponding to a transform in either the streamwise or spanwise direction, the asterisk denotes the complex conjugate, and $k_{x}$ and $k_{y}$ are the associated streamwise and spanwise components of the wave vector, respectively. In order to allow for a direct comparability of these spectra with their corresponding term in the kinetic energy or buoyancy variance equation – see equations (26) and (27), respectively, in appendix A – via Parseval’s theorem, the coefficient $C$ depends on the particular quantities $\Phi_{1}$ and $\Phi_{2}$: $C=1/2$ for $\Phi_{1}=\Phi_{2}\in\left\{u_{x},u_{y},u_{z},b\right\}$, $C=\mathrm{Ri}_{0}$ for $\Phi_{1}=u_{z}\wedge\Phi_{2}=b$ (or vice versa), $C=2/\mathrm{Re}_{0}$ for $\Phi_{1}=\Phi_{2}=\bm{S}$, or $C=1/\left(\mathrm{Re}_{0}\mathrm{Pr}\right)$ for $\Phi_{1}=\Phi_{2}=\nabla b$. We summarise the key results in figure 7.
Figure 7: Magnitude quantification of emergent structures. (a) Emergent flow structures offer pronounced spectral peaks $\hat{\lambda}$ in the streamwise directions for all variables $\left\{u_{x},u_{y},u_{z},b\right\}$. In contrast, (b) only the streamwise velocity $u_{x}$ exhibits a pronounced peak in spanwise direction. (c) A systematic comparison of characteristic extensions of emergent flow structures along the streamwise, spanwise, and vertical directions highlights a strong anisotropy in large-scale dynamics. Note that while the coloured lines in panels (a, b) correspond to our largest square domain with $L_{\mathrm{h}}=512$, grey lines are extracted from even more extended but non-square domains ($L_{x}=2048$ and $L_{y}=512$ or $L_{y}=2048$ and $L_{x}=512$). Moreover, as the dash-dotted line $L_{x}=L_{y}$ in panel (c) demonstrates that flow structures may clearly be limited or affected by horizontally insufficiently extended domains, only horizontal extents $L_{\mathrm{h}}\gtrsim 256$ are large enough to resolve the most extended flow structures in the horizontal direction.
In panel (a), we plot the streamwise energy spectra of $\left\{u_{x},u_{y},u_{z},b\right\}$, subject to appropriate spatio-temporal averages. Remarkably, we find that all these flow fields exhibit pronounced spectral peaks at streamwise wave numbers $\hat{k}_{x}\sim\mathcal{O}\left(10^{-2}\right)$. This implies that a certain (narrow range of) streamwise wavelength(s) $\hat{\lambda}_{x}=2\pi/\hat{k}_{x}$ is particularly energetic. In other words, these pronounced spectral peaks establish the existence of flow structures with a characteristic streamwise length scale $\Lambda_{x}\sim\mathcal{O}\left(10^{2}\right)$ in all of these flow fields, thus implying that there is a preferred length scale for the emergent streamwise self-organisation of the large-scale dynamics. We propose a potential dynamical origin of this characteristic scale $\Lambda_{x}$ in section 4.
As shown in panel (b), this emergent property of the streamwise spectra is in clear contrast to the behaviour of the spanwise spectra. On the one hand, we find a similarly pronounced spectral peak only for the $x$-component of the velocity field in the spanwise direction. Together with $\hat{k}_{y}\sim\mathcal{O}\left(10^{-1}\right)$, this implies again the existence of a characteristic spanwise length scale $\Lambda_{y}\sim\mathcal{O}\left(10^{1}\right)$. On the other hand, the other flow fields do not exhibit such a peak but rather flatten out at small $k_{y}$, demonstrating that there is no preferred characteristic length scale for the spanwise self-organisation of the large-scale dynamics in $\left\{u_{y},u_{z},b\right\}$. Note that even though $E_{u_{y}u_{y}}(k_{y,\textrm{min}})<E_{u_{y}u_{y}}(\hat{k}_{y})$, this difference is smaller than a factor of two and so we do not consider the associated maximum to be a ‘pronounced’ spectral peak.
It is natural to ask whether there could be additional spectral peaks at even smaller wave numbers, i.e., whether our largest square domains of $L_{\mathrm{h}}=512$ might still be too small to capture additional emergent dynamics at even larger length scales. For this reason, we have conducted two additional simulations which increase either the streamwise or the spanwise extent of the domain by another factor of four. This results in $L_{x}\times L_{y}=2048\times 512$ or $512\times 2048$, respectively. We include the associated energy spectra in figure 7 (a, b) as grey solid lines. We find no evidence of such additional spectral peaks. Furthermore, as both the location and amplitude of the peaks from these spectra coincide with the ones from our largest square domain, these spectral peaks are shown to be characteristic of this particular flow and (crucially) independent of the horizontal extent of the domain. This implies that the large-scale dynamics is (at such large $L_{x}$ and $L_{y}$) indeed governed by mechanisms intrinsic to the flow. This is supported by appendix C, where we show that the energy spectra derived from smaller (yet sufficiently large) domains are also shown to converge with the present ones from the largest domains.
streamwise $\hat{\lambda}$ spanwise $\hat{\lambda}$ vertical extent
$L_{x}\times L_{y}\thickspace\thickspace$ $E_{u_{x}u_{x}}$ $E_{u_{y}u_{y}}$ $E_{u_{z}u_{z}}$ $E_{bb}$ $E_{u_{z}b}$ $E_{u_{x}u_{x}}$ $2\times d$ $2\times\delta$
$16^{2}$ 00 $016\pm\hphantom{0}$ 0 $16\pm\hphantom{0}0$ $16\pm\hphantom{0}0$ $016\pm\hphantom{0}$ 0 $16\pm\hphantom{0}0$ $15\pm\hphantom{0}3$ 05.3±0.3plus-or-minus5.30.35.3$\pm$ 0.35.3 ± 0.3 05.6±0.3plus-or-minus5.60.35.6$\pm$ 0.35.6 ± 0.3
$32^{2}$ 00 $032\pm\hphantom{0}$ 0 $32\pm\hphantom{0}0$ $32\pm\hphantom{0}0$ $032\pm\hphantom{0}$ 0 $32\pm\hphantom{0}0$ $19\pm\hphantom{0}6$ $10.7\pm 0.7$ $11.1\pm 0.7$
$64^{2}$ 00 $064\pm\hphantom{0}$ 0 $64\pm\hphantom{0}0$ $64\pm\hphantom{0}2$ $064\pm\hphantom{0}$ 0 $64\pm\hphantom{0}0$ $33\pm 14$ $15.1\pm 0.9$ $15.5\pm 0.7$
[…]
$128^{2}$ 00 $128\pm\hphantom{0}0$ $89\pm 32$ $77\pm 27$ $108\pm 30$ $89\pm 32$ $52\pm 24$ $16.1\pm 0.8$ $16.8\pm 0.9$
$256^{2}$ 00 $117\pm 18$ $84\pm 13$ $74\pm 12$ $92\pm 20$ $80\pm 12$ $52\pm 27$ $15.8\pm 0.4$ $16.4\pm 0.3$
$512^{2}$ 00 $108\pm 19$ $82\pm 10$ $76\pm\hphantom{0}9$ 089±11plus-or-minus891189$\pm$ 1189 ± 11 $79\pm\hphantom{0}8$ $49\pm 15$ $15.5\pm 0.2$ $16.1\pm 0.1$
$2048\times 512$ $111\pm 17$ $86\pm 12$ $77\pm 10$ 096±12plus-or-minus961296$\pm$ 1296 ± 12 $91\pm 10$ $46\pm 10$ $15.6\pm 0.1$ $16.2\pm 0.1$
$512\times 2048$ $118\pm 13$ $83\pm 10$ $75\pm\hphantom{0}8$ 095±11plus-or-minus951195$\pm$ 1195 ± 11 $86\pm 12$ $45\pm 15$ $15.5\pm 0.1$ $16.2\pm 0.1$
Table 2: Emergent dynamical scales. This table quantifies the streamwise and spanwise or vertical extent of emergent large-scale dynamics via the wavelength associated with the spectral peak, the shear layer half-depth or the density interface half-depth. Values of the temporal mean and standard deviation are listed, the latter of which might be significant due to the discrete nature of wave numbers. Unreliable values, i.e. when the domains are insufficiently extended, are displayed in grey.
Quantifying characteristic length scales associated with the emergent large-scale dynamics, we extract the wavelengths corresponding to these pronounced spectral peaks (or these spectra’s global maxima) for each simulation and summarise them in table 2. Additionally, figure 7 (c) visualises this dependence on $L_{\mathrm{h}}$. For small $L_{\mathrm{h}}$, as highlighted by the dash-dotted line, these extracted horizontal length scales $\Lambda_{\textrm{x, y}}$ are clearly biased by the domain size. However, this changes once $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$ with the critical horizontal aspect ratio $L_{\mathrm{h,crit}}$ depending on the solution field (and, thus, the associated final characteristic length scale). While most of the peaks are already properly represented given $L_{\mathrm{h,crit}}=96$, some of them – such as $\hat{\lambda}_{x}\left(E_{u_{x}u_{x}}\right)$ – may require $L_{\mathrm{h}}=256$ to be fully resolved. Hence, a full convergence of the characteristic length scales associated with the large-scale dynamics is reached for $L_{\mathrm{h}}\gtrsim 256$ only.
In addition to the characteristic streamwise and spanwise length scales associated with the large-scale dynamics, we include their characteristic vertical extent via the (total) shear layer depth in figure 7 (c). Note that since $d\approx\delta$ or $R\approx 1$, these data points equivalently show the (total) buoyancy interface depth. Comparing the characteristic length scales in the three spatial directions (as denoted by different symbol types), there is a clearly apparent anisotropy of the large-scale dynamics. While vertical scales $\Lambda_{\textrm{z}}\approx 16$ (circles) and spanwise scales $\Lambda_{\textrm{y}}\approx 50$ (upright triangles), streamwise scales (sideways triangles) may extend up to $\Lambda_{\textrm{x}}\approx 115$. Consequently, we find a hierarchy $\Lambda_{\textrm{z}}<\Lambda_{\textrm{y}}<\Lambda_{\textrm{x}}$ that spans across one order of magnitude.
Having extracted these characteristic scales of large-scale dynamics, there are two particularly striking observations. Firstly, while $u_{z}$ and $b$ are related to the vertical transport of buoyancy via equation (16), their individual scales $\hat{\lambda}_{x}$ in $E_{u_{z}u_{z}}$ and $E_{bb}$ differ significantly. Secondly, the depth of the mixing zone, representing the characteristic vertical extent $\Lambda_{\textrm{z}}$, actually converges at smaller domain extents $L_{\mathrm{h}}$ compared with $\Lambda_{\textrm{x}}$ from $E_{bb}$. These two observations suggest the natural question as to why or how can the vertical characteristics (in both the velocity and buoyancy field) converge before the domain is sufficiently large to allow the separate flow fields to converge horizontally?
Figure 8: Buoyancy exchange and dissipation. (a) The buoyancy flux (i.e. the vertical buoyancy advection) across the midplane – visualised via $B\left(x,y,z=0,t=t_{\textrm{evo}}\right)$ for $L_{\mathrm{h}}=512$ – generally stabilises the configuration. (b) The associated streamwise co-spectrum establishes the existence of a characteristic corresponding scale via a pronounced peak, which is in contrast to the spanwise direction. (c) Most of the dissipation, however, is associated with smaller scales similar to the mixing zone depth. In panels (b, c), coloured and grey spectra correspond – similar to figure 7 (a, b) – to the largest square and non-square domains, respectively.
To answer this question, we visualise the buoyancy flux at the midplane in figure 8 (a). Remembering that $B<0$ when $u_{z}$ and $b$ have opposite signs – i.e., dense parcels are moving upwards or light parcels are moving downwards –, consistently with figure 4 (d), it is reasonable to expect that the field shown in figure 8 (a) is negative on average. Despite its complexity, this visualisation of $B$ suggests again the presence of some regularity. While this is confirmed via the associated streamwise co-spectrum, as shown in panel (b), there is no such peak in the spanwise direction. We remark that a peak in this co-spectrum indicates the presence of a scale at which $u_{z}$ and $b$ interact most strongly, or, in other words, a scale at which vertical velocity and buoyancy are strongly correlated (or anti-correlated). This interaction scale $\hat{\lambda}_{x}\left(E_{u_{z}b}\right)\approx 85$ with $\hat{\lambda}_{x}\left(E_{u_{z}u_{z}}\right)<\hat{\lambda}_{x}\left(E_{u_{z}b}\right)<\hat{\lambda}_{x}\left(E_{bb}\right)$. Hence, this peak (at least partially) explains the convergence of the vertical stirring and mixing for $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}\gtrsim\hat{\lambda}_{x}\left(E_{u_{z}b}\right)$ despite other parts of the large-scale dynamics not yet being fully resolved.
4 Discussion and perspective
Aiming to improve our fundamental understanding of the dynamics of forced stratified shear flows, motivated by their relevance to many geophysical flows, we have conducted a series of direct numerical simulations in computational domains of different sizes. Although our (forced) flow configuration is prone to primary Kelvin-Helmholtz instability, our configuration additionally and importantly includes a continuous forcing which tends to restore the initial profiles and ensures that the ensuing turbulence can be sustained for arbitrarily long times. Since it is this (at least quasi-) continuous forcing which connects to the motivating geophysical flows, we largely disregard the initial transient but rather focus on the statistically stationary dynamics during later time.
By varying the horizontal extent of the numerical domain from $16$ to $512$ times the initial shear layer (half-) depth $d_{0}$, we have demonstrated the emergence of strongly anisotropic large-scale dynamics. Provided a sufficiently large computational domain, the emergent characteristic length scales are $\Lambda_{z}\approx 16$ in the vertical direction, $\Lambda_{y}\approx 50$ in the spanwise direction, and up to $\Lambda_{x}\approx 115$ in the streamwise direction. In other words, we find an anisotropic hierarchy $\Lambda_{z}<\Lambda_{y}<\Lambda_{x}$ that spans across one order of magnitude. Crucially, the associated flow structures as well as their induced stirring, dissipation and irreversible mixing are remarkably sensitive to the confinement and start to become independent of the computational domain for extended domains of $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$ only. This convergence can be associated with the eventual resolution of the vertical buoyancy flux, which exhibits a peak characteristic streamwise scale $\Lambda_{x}\approx 85$. This scale represents approximately five times the final (total) depth of the turbulent mixing zone. However, it is important to appreciate that, as we have only considered one particular combination of control parameters ($\mathrm{Pr}$, $\mathrm{Re}_{0}$, $\mathrm{Ri}_{0}$, and $t_{\textrm{r}}$), any observed relationship between the various length scales (particularly with reference to the final depth of the turbulent mixing zone) should be treated with caution.
As shown in section 3.2, for our particular choice of parameters at least, the depth of the turbulent mixing zone increases as the domain becomes horizontally more extended. At one level, this can be explained by an increased number of degrees of freedom of the underlying dynamical system, leading to a higher potential complexity, and thus more vigorous turbulence, causing both a stronger macroscopic stirring and a stronger microscopic mixing. However, there also appears to be a ‘tuning’ of the overall strength of the stratification towards an equilibrium state. Mixing actually increases the ‘strength’ of the stratification (as quantified by either the bulk Richardson number $\mathrm{Ri}$ or appropriate averages of the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$) through deepening the shear layer depth $\langle d\rangle_{t}$ and the buoyancy interface thickness $\langle\delta\rangle_{t}$, see again figure 4. For the turbulence to remain sustained, this deepening cannot continue without limit since turbulence cannot be sustained when the stratification becomes too strong. The particular observation that the (mixing zone) average of $\mathrm{Ri}_{\textrm{g}}$ is approximately $0.15$ is highly reminiscent of the results reported by Portwood et al. (2019). They used a control strategy (effectively through modulating gravity) to identify equilibrium turbulent states in linearly stratified flows driven by constant vertical shear. Although the flow considered here is different, the key dynamics appears to be similar: the forced flow adjusts until the stratification is as strong as possible to still allow for vigorous turbulence which is able to stir and mix the buoyancy field essentially as a passive scalar. This is underlined by the fact that $\langle d\rangle_{t}\approx\langle\delta\rangle_{t}$, and so the mixing of buoyancy is effectively slaved to the mixing of momentum by the turbulence.
Expecting dissipation to be dominated by the smallest scales in the flow, it might be surprising to see that dissipation (and mixing) statistics only converge for $L_{\mathrm{h,crit}}\gg\eta_{\textrm{K}}$ (with the Kolmogorov scale $\eta_{\textrm{K}}\sim\mathcal{O}\left(10^{-1}\right)$, see appendix B). Questioning this potential expectation, we plot the energy spectra associated with the dissipation rates of kinetic energy and buoyancy variance in figure 8 (c). Note that we show them in a pre-multiplied form (Krug et al., 2020) in order to highlight visually those wave numbers that cause most of the variance. We find the peaks in these spectra to be located at $\hat{k}\approx\left(3\dots 7\right)\times 10^{-1}$ or $\hat{\lambda}\approx 9\dots 18$, i.e., being again actually associated with the buoyancy interface depth or equivalently the turbulent mixing zone (due to the above-mentioned slaving). As we have shown in figure 8 (b) and explained in section 3.4, a convergence of this interface depth depends on resolving the streamwise spectral peak of the vertical buoyancy flux and requires $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$, explaining the convergence of the dissipation fields only in extended domains.
Based on our spectral analysis from section 3.4, we have identified characteristic streamwise scales $\Lambda_{x}\approx 75\dots 115$ associated with the large-scale dynamics. We believe that this characteristic length scale can once again be interpreted in terms of the (sustained) emergent shear layer half-depth $\langle d\rangle_{t}\approx 8$ (which is measured in units of $d_{0}$). We remark that, although the imposed forcing is in principle designed to relax the flow back to the initial shear layer half-depth $d_{0}$ (and buoyancy interface half-depth $\delta_{0}$), the (above described) sustained turbulence results in or ensures this significantly deeper shear layer and buoyancy interface. It is well-known (see for example Scinocca & Ford (2000)) that the most unstable mode of KHI has a characteristic wavelength $\lambda_{\textrm{KHI}}$ of approximately fifteen times the shear layer (half-) depth. Here, at late times when $\langle d\rangle_{t}\approx 8$, this implies that $\lambda_{\textrm{KHI}}\sim 120$ which is very similar to the emergent streamwise length scales, particularly those associated with the streamwise velocity field as listed in table 2. However, this preferred instability scale is not able to roll up completely into coherent Kelvin-Helmholtz billows, and in particular, the turbulence certainly disrupts any possibility of subharmonic pairing or merging occuring, analogously to the disruption arguments put forward by Mashayek & Peltier (2013). Furthermore, our observation that spanwise scales are both significantly smaller and significantly harder to identify is consistent with the complete lack of billow coherence, as the knot/tube/defect structure is observed experimentally to require billows of at least $6003-5\lambda_{\textrm{KHI}}\sim 360-$ 600 spanwise extent, which is an order of magnitude larger than the observed spanwise scale $\Lambda_{y}\approx 50$.
In summary, our research leads to two main conclusions. Firstly, it adds to an increasing body of evidence that forced stratified shear flows ‘tune’ (in particular in the vertical direction) to a state with typical values of $\mathrm{Ri}_{\textrm{g}}\lesssim 0.2$. Such a state appears to allow both sustained turbulence and non-trivial buoyancy flux and attendant irreversible mixing in a fundamentally ‘weakly stratified’ regime. In this regime, buoyancy is slaved to the velocity field, and behaves, at least in some sense as a passive scalar, with relatively ‘inefficient’ mixing, in that bulk or average flux coefficients are bounded above by $0.2$, although occasional extreme events with substantially higher instantaneous local flux coefficients (see figure 5) can occur.
Secondly, it suggests the emergence of some imprint of the dominant (and inherently linear) shear-driven instability scale even in a turbulent flow. For numerical simulations to yield converged and reliable statistics therefore, domains should be scaled using the ‘tuned’ emergent vertical depth of the turbulent shear layer, which (during this tuning process) may be significantly larger than its initial value.
Clearly, these conclusions need to be tested for other choices of the key parameters. It would be very instructive to investigate the sensitivity of the ‘tuning’ process to the choice of the initial bulk Richardson number and Reynolds number. Moreover, an application to the ocean would clearly require the investigation of the sensitivity of the flow dynamics to the choice of $Pr$, as thermally stratified water has $Pr\sim\mathcal{O}\left(10\right)$. As we have seen, the emergent scales are both clearly related to the turbulent processes but also enormously larger than the smallest scales of that turbulence. A particular issue of interest is to consider larger initial $\mathrm{Re}_{0}$ that would be unstable right from the beginning of the simulation, avoiding the (potentially confusing) phenomenon of the shear layer effectively doubling in depth before the onset of the primary instability. If our conclusions prove to be robust and generic for other parameter choices, they have major implications for idealized computational studies, as they strongly suggest that to capture key characteristics of sheared turbulence in such ‘weak’ stratification, significantly extended computational domains are required.
There are also interesting implications for the geophysical application of such idealized studies of shear-driven stratified turbulence which completely ignore any effects of rotation. As our study suggests that streamwise scales can emerge that are $\mathcal{O}\left(50\right)$ times larger than the shear layer (half-depth) scale at initial onset of KHI, it is entirely plausible that such larger scales will at least be affected somewhat by rotational effects. As noted in the review of Taylor & Thompson (2023), ‘submesoscale motions’ with Rossby number $\mathrm{Ro}:=U/(fL)\sim O(1)$, where $f$ is the Coriolis parameter and $U$ and $L$ are characteristic velocity and length scales, are flows where ‘the Coriolis acceleration is important, but it does not constrain the motion’. In the world’s oceans, they ‘define submesoscales as dynamical features with horizontal scales between approximately 200 m and 20 km’. It is entirely plausible that geophysical flows such as estuarine outflows could have initial shear layer half-depths of the order of 4-10 m, and that KHI would onset immediately associated with that shear layer depth as observed for example by Holleman et al. (2016). In such flows, the emergent streamwise length scales may well experience non-trivial rotational effects. Therefore, it seems an interesting important question to investigate to what extent the emergent flow properties we have identified might be affected by larger-scale rotation.
[…]
[Acknowledgements] P.P.V. (i) thanks Roshan J. Samuel (TU Ilmenau, Germany) for early help with NekRS and (ii) gratefully acknowledges the support of Pembroke College, Cambridge, through a postdoctoral research associateship.
[…]
[Funding] P.P.V. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within Walter Benjamin Programme 532721742. The authors gratefully acknowledge the Gauss Centre for Supercomuting e.V. (www.gauss-centre.eu) for funding this work by providing computing resources through the John von Neumann Institute for Computing (NIC) on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC) within project nonbou.
[…]
[Declaration of interests] The authors report no conflict of interest.
[…]
[Author ORCIDs] P. P. Vieweg, https://orcid.org/$0$-0001-7628-9902
C. P. Caulfield, https://orcid.org/$0$-0002-3170-9480
[…]
[Author contributions] Both authors designed the study. P.P.V. performed the numerical simulations, processed the generated data, and drafted the manuscript. Both authors contributed equally in discussing the data and finalising the paper.
Appendix A Kinetic energy and buoyancy variance equations
All of our simulations are based on equations (3) and (4) as the evolution equations for the momentum and buoyancy, respectively. However, some quantities used in the main text are actually related to the evolution equations of kinetic energy $\bm{u}^{2}/2$ or buoyancy variance $b^{2}/2$ instead. These two (non-dimensional) equations are
$\frac{\partial}{\partial t}\frac{\bm{u}^{2}}{2}$ $=-\nabla\cdot\bm{J}_{u}-\varepsilon_{u}+F_{u}+B,$ (26)
$\frac{\partial}{\partial t}\frac{b^{2}}{2}$ $=-\nabla\cdot\bm{J}_{b}-\varepsilon_{b}+F_{b}$ (27)
and follow from taking the dot product of equation (3) with $\bm{u}$ or multiplying (4) by $b$. Both equations have a similar structure.
The dissipation rates $\varepsilon_{\Phi}$ have already been introduced in the main text in equations (17) and (18), whereas the vertical buoyancy advection or buoyancy flux $B$ has been introduced in equation (16). In addition, the $j$-th components of the flux terms $\bm{J}_{\Phi}$ are given by
$J_{u,j}$ $:=u_{j}\frac{u_{i}^{2}}{2}+u_{j}p-\frac{1}{\mathrm{Re}_{0}}\left[\frac{\partial}{\partial x_{j}}\frac{u_{i}^{2}}{2}+\frac{\partial}{\partial x_{i}}\left(u_{i}u_{j}\right)\right],$ (28)
$J_{b,j}$ $:=u_{j}\frac{b^{2}}{2}-\frac{1}{\mathrm{Re}_{0}\mathrm{Pr}}\frac{\partial}{\partial x_{j}}\frac{b^{2}}{2}$ (29)
whereas our additional source terms $F_{\Phi}$ are given by
$F_{u}:=u_{x}f_{u}\qquad\textrm{and}\qquad F_{b}:=bf_{b}.$ (30)
Note that we apply the Einstein summation convention in equations (28) and (29) for improved clarity. The volumetric forcings $f_{\Phi}$ have been defined in the main text in equations (6) and (7).
Equation (27) represents the evolution equation for the unscaled buoyancy variance. The evolution equation for the scaled buoyancy variance follows simply from a subsequent multiplication of this equation by $\mathrm{Ri}_{0}^{2}/\langle N^{2}\rangle_{A,t}$, see also again equations (18), (20), and (21).
Appendix B Resolving shear flows efficiently using spectral element methods
Direct numerical simulations aim to resolve all dynamically relevant scales, from the domain size down to the Kolmogorov scale or Batchelor scale. In our non-dimensional description, see again section 2.1 and equation (1), these smallest scales are locally given by
$\eta_{\textrm{K}}:=\left(\frac{1}{\mathrm{Re}_{0}^{3}\thickspace\varepsilon_{u}}\right)^{1/4}\qquad\textrm{and}\qquad\eta_{\textrm{B}}:=\frac{\eta_{\textrm{K}}}{\sqrt{\mathrm{Pr}}},$ (31)
and both depend on the kinetic energy dissipation rate $\varepsilon_{u}$ as defined by equation (17).
Figure 9: Resolving shear flows. (a) Shear flows exhibit highly non-uniform profiles of dissipation $\varepsilon_{u}$ and, thus, the smallest dynamical scales $\eta_{\textrm{K}}$. (b) Spectral element methods allow for the adjustment of the height of spectral elements $h_{\mathrm{se}}$. The local resolution $dz$ follows from a subsequent spectral expansion of polynomial order $N$ within each element. Hence, adjusting $h_{\mathrm{se}}$ enables (c) the very efficient resolution of shear flows. Here, just like in figure 2, $L_{\mathrm{h}}=128$.
As already presented in figure 3, shear flows – as visualised also in figure 2 – offer highly non-uniform spatial distributions of dissipation, including this kinetic energy dissipation rate. For example, figure 9 re-plots this $\langle\varepsilon_{u}\rangle_{A,t}\left(z\right)$ in panel (a). The dissipation is strongest at midplane but vanishes sufficiently far away. We compute the associated local Kolmogorov scale $\langle\eta_{\textrm{K}}\rangle_{A,t}$ from $\langle\varepsilon_{u}\rangle_{A,t}$ to obtain a proxy for the (on average) smallest local dynamical scale.
The local (vertical) resolution $dz$ of spectral element methods is determined by the interplay between the size (here: height) of each spectral element $h_{\textrm{se}}$ and the polynomial order $N$ (Deville et al., 2002; Vieweg, 2023). It has empirically been shown (Scheel et al., 2013) that spectral element methods offer smooth dissipation fields, even at the spectral element boundaries, when the (refined) Grötzbach criterion (Scheel et al., 2013; Vieweg, 2023)
$\frac{dz}{\langle\eta_{\textrm{K}}\rangle_{A,t}}\lesssim\frac{\pi}{2}\quad\textrm{for }\thickspace\mathrm{Pr}\leq 1\qquad\textrm{or}\qquad\frac{dz}{\langle\eta_{\textrm{B}}\rangle_{A,t}}\lesssim\frac{\pi}{2}\quad\textrm{for }\thickspace\mathrm{Pr}\geq 1$ (32)
is satisfied. It is common to use this as a criterion for the spatial resolution.
Hence, as shown in panel (b), we adjust $h_{\textrm{se}}\left(z\right)$ to mimic the observed trends in $\langle\eta_{\textrm{K}}\rangle_{A,t}\left(z\right)$. After the spectral expansion within each spectral element, the local $dz$ follows a similar trend. As highlighted eventually by panel (c), this procedure of adjusting the local spectral element height enables the successful and efficient resolution of shear flows.
Finally, we remark that we derived $h_{\textrm{se}}\left(z\right)$ from preliminary tests and use the same distribution for all our production simulations. The horizontal width of spectral elements $w_{\textrm{se}}$ is constant and similar to $h_{\textrm{se}}\left(z=0\right)$. A comparison of time-averaged vertical profiles between our $L_{z}=48$ domain and a vertically even more extended domain of $L_{z}=56$ has confirmed their convergence during these preliminary tests.
Appendix C Convergence of horizontal energy spectra
Figure 10: Convergence of energy spectra. Both the (a – d) streamwise and (e – h) spanwise energy spectra converge for $L_{\mathrm{h}}\gtrsim 96$ to the accessible parts of the spectra from even larger domains. Thin black lines are extracted from the most extended but non-square domains ($L_{x}=2048$ and $L_{y}=512$ or vice versa).
In section 3.4, we have introduced Fourier energy spectra and explained that the location of their peak serves as a measure to quantify a length scale which is characteristic of the large-scale dynamics. That analysis has been conducted for our largest square domain at $L_{\mathrm{h}}=512$ and underlined by even larger but non-square domains, see particularly figure 7.
In this appendix, via figure 10, we additionally contrast the Fourier energy spectra from all our simulations with different $L_{\mathrm{h}}$. First and foremost, we confirm that these energy spectra can indeed overlap given a sufficient $L_{\mathrm{h}}$. We find that the energy spectra obtained from smaller domains $L_{\mathrm{h}}<\Gamma_{\textrm{h, max}}$ have indeed converged (to those from our largest domains) once $L_{\mathrm{h}}\gtrsim 96$. In contrast, smaller domains offer deviating spectra, especially along the spanwise direction as can be seen from panels (e – h).
Let us stress that convergence of spectra means that these spectra overlap at mutually accessible wave numbers. Remember that this accessible range of wave numbers is limited by the smallest positive wave number $k_{\textrm{min}}=2\pi/L_{\mathrm{h}}$ due to the domain size $L_{\mathrm{h}}$. So for instance, although the spectrum of $E_{u_{x}u_{x}}$ from both $L_{\mathrm{h}}=\left\{96,512\right\}$ overlaps for $k\geq k_{\textrm{min}}\left(L_{\mathrm{h}}=96\right)\approx 0.07$, there is no spectrum accessible at $k<k_{\textrm{min}}\left(L_{\mathrm{h}}=96\right)$ for a domain of $L_{\mathrm{h}}=96$.
This convergence has important implications. Even if a domain is too small to resolve entirely the spectral peaks $\hat{\lambda}$ forming at extreme $L_{\mathrm{h}}$, see again table 2, the accessible part of the spectrum has still converged. In other words, the dynamics at smaller scales is then no longer affected by the missing part of the spectrum or dynamics.